OpenStudy (anonymous):

help?!? Two birds sit at the top of two different trees 17.8 feet away from one another. The distance between the first bird and a bird watcher on the ground is 37.4 feet. What is the angle measure, or angle of depression, between this bird and the bird watcher?(will give medal)

OpenStudy (anonymous):

25.5° 64.5° 28.4° 61.6°

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f38a277e4b0fc0c1a0dd52f It is already solved please look it here

OpenStudy (anonymous):

@SolomonZelman can u help

OpenStudy (solomonzelman):

Do you know what \[a^2+b^2=C^2\] is?

OpenStudy (solomonzelman):

I didn't mean a capitol c.

OpenStudy (anonymous):

yes Pythagorean therom

OpenStudy (anonymous):

Pythogaras thm

OpenStudy (solomonzelman):

Plug in your numbers.

OpenStudy (solomonzelman):

\[37.4^2-17.8^2=x^2\]

OpenStudy (solomonzelman):

Solve for x.

OpenStudy (anonymous):

1398-316=1082

OpenStudy (anonymous):

@SolomonZelman is it 1398-316=1082

OpenStudy (solomonzelman):

yes, but please try to write at least 2 decimal places. 1398.76-316.84=1081.92 now take a square root of this, or you need a precise simplified answer?

OpenStudy (solomonzelman):

 exact answer, or just take a square root?